Suppose the labor market has the following statistics: Unemployed=1million, Employed=4million, Adult population= 7million. (a) Suppose of...

  1.   Suppose the labor market has the following statistics: Unemployed=1million, Employed=4million, Adult population= 7million.

(a) Suppose of the 1million unemployed, 80,000 of them find jobs every month; of the 4million employed,160,000 lose their jobs every month.

(i) Calculate and interpret the steady state rate of unemployment

(ii) Give an example of a policy that would reduce the rate of job separation.

(iii) Suppose this policy reduces the number of people losing their jobs from 160,000 to 120,000. Calculate and explain how this will affect the steady state unemployment rate
Answer #1

i) Notations:

L =Number of workers in labor force = (E+U= 5 million)

E = Number of employed workers = 4 millions

U = Number of unemployed = 1 million

U/L = unemployment rate = 1million/5 million=0.2 or 20% and if we consider the total Labour force (L) as 7 million, the unemployment rate will be =1million/7 million=0.14 or 14%.

If we assume

L is exogenously fixed as 5 million.

During any given month, s = rate of job separations, the fraction of employed workers that become separated from their jobs = 1,60,000= 0.004 and

f = rate of job finding, fraction of unemployed workers that find jobs = 80000=0.008

s and f are exogenous

The steady-state condition is: sE = fU

Finding the “equilibrium” U rate

f U = sE

= s(L –U )

= s*L – s*U

Solve for U/L: (f + s)U = s*L

So U/L = s/s+f= 0.004/0.004+0.008=0.33 or 33%

ii) A policy will reduce the natural rate of unemployment only if it lowers s or increases f

iii) Suppose this policy reduces the number of people losing their jobs from 160,000 to 120,000

the s = rate of job separations will be 0.003

so the steady state rate of Unemployment =U/L=0.003/0.003+0.008=0.27 or 27%